# Dreaming of Many Worlds: Learning Contextual World Models Aids Zero-Shot Generalization

### v1

@@ Abstract
Zero-shot generalization (ZSG) to unseen dynamics is a major challenge for creating generally capable embodied agents. To address the broader challenge, we start with the simpler setting of contextual reinforcement learning (cRL), assuming observability of the context values that parameterize the variation in the system’s dynamics, such as the mass or dimensions of a robot, without making further simplifying assumptions about the observability of the Markovian state. Toward the goal of ZSG to unseen variation in context, we propose the contextual recurrent state-space model (cRSSM), which introduces changes to the world model of the Dreamer (v3) (Hafner et al., 2023). This allows the world model to incorporate context for inferring latent Markovian states from the observations and modeling the latent dynamics. Our experiments show that such systematic incorporation of the context improves the ZSG of the policies trained on the “dreams” of the world model. We further find qualitatively that our approach allows Dreamer to disentangle the latent state from context, allowing it to extrapolate its dreams to the many worlds of unseen contexts. The code for all our experiments is available at https://github.com/sai-prasanna/dreaming_of_many_worlds.
@@ 1 Introduction
Here, we propose to use the contextual reinforcement learning paradigm (Kirk et al., 2023; Benjamins et al., 2023) to study Dreamers’ learning capabilities within and across many worlds. To this end, we assume that we have access to privileged information about how the transition dynamics of the underlying Markov decision process (MDP) is parameterized, i.e., the context(Hallak et al., 2015). Following Benjamins et al. (2023), we assume that context defines some physical properties, such as gravity or the mass of a load, that a robot might measure while trying to solve a given task.
@@
We provide the first principled study in understanding Dreamer’s generalization capabilities for in- as well as out-of-distribution (OOD) tasks;
We propose a novel approach for conditioning the Dreamer architecture on context and show how it improves Dreamer’s zero-shot generalization ability and
@@ 2 Related Work
Meta-reinforcement learning (meta-RL)111For a recent survey we refer to Beck et al. (2023). has been proposed as a promising approach to address the challenge of generalization in RL. Meta-RL aims to learn an RL agent that can adapt to new tasks in a sample-efficient manner. Meta-RL algorithms (see, e.g., Duan et al., 2016; Wang et al., 2017; Nagabandi et al., 2019; Rakelly et al., 2019; Melo, 2022; Wen et al., 2023) are designed to quickly adapt to new and unseen settings with limited access to new experiences (i.e., few-shot adaptation) generated by the RL agent. In contrast, our work focuses on zero-shot generalization (ZSG) for RL (Kirk et al., 2023), where we aim to learn policies that are capable of zero-shot adaptation to new settings without assuming access to further training or the reward signal.
@@
Empirically, MBRL algorithms, such as Dreamer (Hafner et al., 2020; 2023) and TD-MPC2 (Hansen et al., 2024) pipelines, achieve state-of-the-art sample efficiency. Based on Dreamer’s success, we build our approach on it and study its zero-shot generalization capabilities before suggesting an improved approach. Dreamer has recently been studied in the meta-RL case for a few-shot generalization (Wen et al., 2023). However, this work still requires many interactions with the target domain for the agent to learn to adapt to a test task successfully.
@@ 3 Background
We often face partial observability of latent Markovian states in real-world RL tasks. Providing context to an agent may help infer such latent Markovian states.
As a motivating example, consider a wheeled robot that has to deliver goods to various locations, encountering varying terrains along the way, including rough terrains that may damage the robot’s wheels and, ultimately, the robot. Now, assume that the robot has a sensor to measure its velocity. If the robot must navigate without damaging itself, it may be useful to estimate the coefficient of friction of the surface on which it moves. This coefficient of friction is a non-stationary latent state variable that changes within an episode.
Here, the mass of the load on the robot (context), the coefficient of friction of the terrain (latent state), and the torque applied to the wheels (action) causally affect the velocity (observation) of the robot. The robot must use its context, observations, and actions to infer the friction. The robot could then use this estimate of the friction to improve its policy and decide whether to apply more or less torque. This example, though oversimplified, gives us an idea of how effective the context might be to learn to infer latent states in partially observable settings and to use the latent state to improve its policy. It motivates the design of our contextual Dreamer agent, which we describe now.
@@ 4.1.1 Contextual Recurrent State-Space Model (cRSSM)
To perform inference of the latent states for this non-linear model, we cannot directly compute the posterior (Hafner et al., 2019). Instead, we learn an encoder . This encoder incorporates context to estimate latent states from observations and actions. Using the encoder, we follow Hafner et al. (2019) in constructing a variational bound on the data log-likelihood. Here, we write the objective for predicting only the observations; a similar derivation applies for the rewards and the prediction of the continuation flag222Continuation flag indicates whether the state is terminal. of the episode . The evidence lower bound (ELBO) obtained using Jensen’s inequality is then
We mainly extend the steps in Hafner et al. 2019 for constructing the lower bound with context. For the derivation, refer to Appendix A. The expectations in this objective can be optimized with gradient ascent on samples drawn from the encoder using the reparameterization trick (Kingma & Welling (2014)).
@@ 4.1.3 Actor-Critic Policy
We largely follow Dreamer(v3) regarding training an actor-critic policy on the imagined trajectories. However, we introduce context into policy learning by conditioning the actor and the critic networks with the context. The actor is optimized to maximize the expected return on the imagined trajectories.
@@ 4.2 Naive Use of Context in Dreamer
As discussed previously, Dreamers ZSG capabilities have not been explored in an observable context setting, nor, to the best of our knowledge, has this setting been explored in MBRL in general. Thus, here, we propose and discuss naive learning variants from and within the contextual setting and contrast them with our proposed cRSSM. The naive variants are then used as baselines in our experiments later.
@@ 4.2.1 Context as an Observation
A commonly adopted approach to incorporate the inferred or true context into an algorithm is to concatenate it with the state or observation (Perez et al., 2020; Biedenkapp et al., 2022; Sodhani et al., 2022; Benjamins et al., 2023). We study applying this approach to vanilla Dreamer. The vanilla Dreamer optimizes all the objectives defined in Section 4.1, but it does so without incorporating context in those objectives. To incorporate context naively, we provide it concatenated with the observation to the stochastic state encoder. where . Note that only the encoder gets the observation (or here the observation with context) as an input. The decoder then has to learn to reconstruct the context as it is part of the observation. The latent dynamics predictor used for imagination does not condition on observations. For a consistent imagination, the RSSM is burdened with retaining the context value (provided as an observation) which got encoded into the latent state, from which imagination begins. Since we only provide the latent state inferred by the encoder to the actor-critic model in this setting, if the context is not retained, then the actor-critic network will also not have access to the context information to learn accurate policies. This could make it hard to generalize OOD. Still, the simplicity of directly using context as part of the observation is an appealing approach, which has led to it being the predominant approach in model-free cRL.
@@ 5 Experiments and Discussion
In this section, we assess the performance of Dreamer in achieving generalization under observed contexts. We compare various context conditioning approaches, following the evaluation protocols outlined in (Kirk et al., 2023). In particular, our findings highlight the effectiveness of our cRSSM method, showcasing quantitative and qualitative results in terms of zero-shot generalization (ZSG).
5.1 Setup
@@
Our experiments leverage the Contextual and Adaptive Reinforcement Learning benchmark (CARL; Benjamins et al., 2023), tailored for our investigation into ZSG. In CARL we pick the following environments and contexts.
@@
Single context variation: We randomly sample context values uniformly for one context dimension in its training range, keeping the other fixed to its default value and vice versa.
@@
In the evaluation protocol, each context could constitute a task of different difficulty. For example, learning to control an agent in a context where it carries lighter loads might be easier than the one with heavier loads. Following Benjamins et al. (2023), to obtain an upper bound of the policy returns, we train expert agents for selected contexts that broadly cover our ranges of training and evaluation contexts. Expert agent performances are the best mean return over episodes among five seeds. This gives an upper bound on the returns achievable if Dreamer is trained in a particular context. We used the best random policy mean return on episodes over five seeds to define a lower bound. Refer to Section C.1 for the performance of expert and random policies in each environment and context value.
@@
| Setting | I | E | I+E | I | E | I+E |
| CartPole | ||||||
| Featurized | Pixel | |||||
| (g d) | 1.000 | 1.000 | - | 1.000 | 0.938 | - |
| (g h) | 1.000 | 1.000 | - | 1.000 | 0.995 | - |
| (g c) | 1.000 | 1.000 | - | 1.000 | 0.997 | - |
| (g cR) | 1.000 | 1.000 | - | 1.000 | 1.000 | - |
| (l d) | 1.000 | 0.995 | - | 0.677 | 0.059 | - |
| (l h) | 1.000 | 0.996 | - | 1.000 | 0.169 | - |
| (l c) | 1.000 | 0.987 | - | 1.000 | 0.210 | - |
| (l cR) | 1.000 | 1.000 | - | 1.000 | 0.374 | - |
| (g+l d) | 1.000 | 0.945 | 0.998 | 0.901 | 0.038 | 0.210 |
| (g+l h) | 1.000 | 0.989 | 1.000 | 1.000 | 0.149 | 0.701 |
| (g+l c) | 1.000 | 0.970 | 1.000 | 1.000 | 0.257 | 0.779 |
| (g+l cR) | 1.000 | 0.997 | 1.000 | 1.000 | 0.334 | 0.826 |
| Walker | ||||||
| Featurized | Pixel | |||||
| (g d) | 0.903 | 0.561 | - | 0.940 | 0.546 | - |
| (g h) | 0.967 | 0.764 | - | 0.945 | 0.708 | - |
| (g c) | 0.966 | 0.769 | - | 0.966 | 0.733 | - |
| (g cR) | 0.985 | 0.806 | - | 0.959 | 0.710 | - |
| (a d) | 0.885 | 0.479 | - | 0.959 | 0.461 | - |
| (a h) | 0.959 | 0.571 | - | 0.947 | 0.571 | - |
| (a c) | 0.926 | 0.597 | - | 0.983 | 0.635 | - |
| (a cR) | 0.998 | 0.674 | - | 0.994 | 0.623 | - |
| (g+a d) | 0.842 | 0.520 | 0.595 | 0.915 | 0.503 | 0.570 |
| (g+a h) | 0.966 | 0.764 | 0.843 | 0.952 | 0.666 | 0.772 |
| (g+a c) | 0.972 | 0.727 | 0.830 | 0.965 | 0.724 | 0.823 |
| (g+a cR) | 0.982 | 0.677 | 0.820 | 0.988 | 0.691 | 0.843 |
@@ 5.2.1 How Effective is Domain Randomization for Dreamer’s ZSG?
To answer this, we compare the two approaches default-context, where the agent is trained on the default context , and the hidden-context, which involves training the agent with domain randomization of contexts.
@@
In summary, domain randomization benefits the ZSG of the Dreamer algorithm, and the improvement is striking for the pronounced pixel modality.
@@ 5.2.2 Does Explicit Context Conditioning Aid ZSG?
Having established the benefits of domain randomization through hidden-context for Dreamer’s ZSG, our focus shifts to evaluating the impact of explicit context conditioning methods, namely cRSSM, our principled way to incorporate context into Dreamer’s world model; and concat-context where we augment the observations with the context.
@@ 5.2.3 Beyond Task Performance, Does Context-Conditioning Impact the Latent States?
To evaluate how the methods use context, we visually investigate Dreamer’s reconstruction of visually observable OOD contexts, which are either encoded into the latent state in concat-context or conditioned separately cRSSM. We choose the visually observable context to be the pole length in Cartpole, evaluating lengths shorter and longer than what it has seen in the training distribution (labeled short and long in Figure 5). A model capable of generating images conditioned on novel context demonstrates a semantic understanding of the context that is well grounded in the image space. In scenarios involving context-conditioned models, we additionally provide counterfactual visual explanations, exploring how the reconstructed pixel observation inferred from the original pixel observation that encodes the factual context () would differ if conditioned on a counterfactual context ().
@@
Our investigation reveals clear evidence of extrapolation capabilities in our proposed principled cRSSM approach compared to the vanilla concat-context strategy. Furthermore, through our visual counterfactual explanations, we observe indications that the latent state in cRSSM appears disentangled from the context, which explains the observed gains in generalization afforded by this approach.
@@ 6 Conclusion and Future Work
We studied zero-shot generalization in Dreamer-style model-based reinforcement learning through the lens of contextual reinforcement learning. We discussed naive ways to incorporate contextual information into the MBRL learning pipeline and formulated the novel cRSSM for Dreamer. Our cRSSM provides a systematic approach to using context in the world modelling objectives under partial observability. Our experiments, using a rigorous evaluation protocol for zero-shot generalization, showed that naive approaches, such as domain randomization improve generalization performance. However, more principled methods such as our cRSSM are required to perform significantly better in-distribution and out-of-distribution. Our study opens the door to future work on zero-shot generalization for MBRL approaches such as Dreamer.

### v2

@@ Abstract
Zero-shot generalization (ZSG) to unseen dynamics is a major challenge for creating generally capable embodied agents. To address the broader challenge, we start with the simpler setting of contextual reinforcement learning (cRL), assuming observability of the context values that parameterize the variation in the system’s dynamics, such as the mass or dimensions of a robot, without making further simplifying assumptions about the observability of the Markovian state. Toward the goal of ZSG to unseen variation in context, we propose the contextual recurrent state-space model (cRSSM), which introduces changes to the world model of Dreamer (v3) (Hafner et al., 2023). This allows the world model to incorporate context for inferring latent Markovian states from the observations and modeling the latent dynamics. Our approach is evaluated on two tasks from the CARL benchmark suite, which is tailored to study contextual RL. Our experiments show that such systematic incorporation of the context improves the ZSG of the policies trained on the “dreams” of the world model. We further find qualitatively that our approach allows Dreamer to disentangle the latent state from context, allowing it to extrapolate its dreams to the many worlds of unseen contexts. The code for all our experiments is available at https://github.com/sai-prasanna/dreaming_of_many_worlds.
@@ 1 Introduction
Here, we propose to use the contextual reinforcement learning paradigm (Kirk et al., 2023; Benjamins et al., 2023) to study Dreamers’ learning capabilities within and across many worlds. To this end, we assume that we have access to privileged information about how the transition dynamics of the underlying Markov decision process (MDP) is parameterized, i.e., the context(Hallak et al., 2015). We use tasks from the Contextual and Adaptive Reinforcement Learning benchmark (CARL; Benjamins et al., 2023) where the context defines some physical properties that affect the dynamics and that an RL agent can observe while trying to solve a given task. Examples of such properties are gravity or the mass of a load that a robot might lift. We assume the context parameters are continuous and hence can be studied meaningfully for interpolation and extrapolation, unlike discrete parameters where generalization to unseen values is not well-defined.
@@
We provide the first principled study in understanding Dreamer’s generalization capabilities for in- as well as out-of-distribution (OOD) tasks on two tasks from CARL;
We propose a novel approach for conditioning the Dreamer architecture on context and show how it improves Dreamer’s zero-shot generalization ability on the given tasks; and
@@ 2 Related Work
Meta-reinforcement learning (meta-RL) (Beck et al., 2023) has been proposed as a promising approach to address the challenge of generalization in RL. Meta-RL aims to learn an RL agent that can adapt to new tasks in a sample-efficient manner. Meta-RL algorithms (see, e.g., Duan et al., 2016; Wang et al., 2017; Nagabandi et al., 2019; Rakelly et al., 2019; Melo, 2022; Wen et al., 2023) are designed to quickly adapt to new and unseen settings with limited access to new experiences (i.e., few-shot adaptation) generated by the RL agent. In contrast, our work focuses on zero-shot generalization (ZSG) for RL (Kirk et al., 2023), where we aim to learn policies that are capable of zero-shot adaptation to new settings without assuming access to further training or the reward signal.
@@
Empirically, MBRL algorithms, such as Dreamer (Hafner et al., 2020; 2023) and TD-MPC2 (Hansen et al., 2024) pipelines, achieve state-of-the-art sample efficiency. While Dreamer’s focus was on achieving a high return in a variety of individual environments, our focus is to study and improve generalization capabilities in contextual variants of environments and we evaluate our work on currently available environments tailored for this. Based on Dreamer’s success, we build our approach on it and study its zero-shot generalization capabilities before suggesting an improved approach. Dreamer has recently been studied in the meta-RL case for a few-shot generalization (Wen et al., 2023). However, this work still requires many interactions with the target domain for the agent to learn to adapt to a test task successfully.
@@ 3 Background
We often face partial observability of latent Markovian states in real-world RL tasks. Providing context to an agent may help infer such latent Markovian states. Generally, context refers to an aspect of the environment or the agent that remains constant for a certain period and affects its behavior. In our work, we focus on context that remains unchanged throughout the entire episode.
As a motivating example, consider a wheeled robot that has to deliver goods to various locations, encountering varying terrains along the way, including rough terrains that may damage the robot’s wheels and, ultimately, the robot. Now, assume that the robot has a sensor to measure its velocity. If the robot must navigate without damaging itself, it may be useful to estimate the coefficient of friction of the surface on which it moves. This coefficient of friction is a non-stationary latent state variable that changes within an episode. Here, the mass of the load on the robot (context), the coefficient of friction of the terrain (latent state), and the torque applied to the wheels (action) causally affect the velocity (observation) of the robot. The robot must use its context, observations, and actions to infer the friction. The robot could then use this estimate of the friction to improve its policy and decide whether to apply more or less torque. This example, though oversimplified, gives us an idea of how effective the context might be to infer latent states and improve the policy in partially observable settings. It motivates the design of our contextual Dreamer agent.
@@ 4.1.1 Contextual Recurrent State-Space Model (cRSSM)
To perform inference of the latent states for this non-linear model, we cannot directly compute the posterior (Hafner et al., 2019). Instead, we learn an encoder . This encoder incorporates context to estimate latent states from observations and actions. Using the encoder, we follow Hafner et al. (2019) in constructing a variational bound on the data log-likelihood. Here, we write the objective for predicting only the observations; a similar derivation applies for the rewards and the prediction of the continuation flag111Continuation flag indicates whether the state is terminal. of the episode . The evidence lower bound (ELBO) obtained using Jensen’s inequality is then
We mainly extend the steps in Hafner et al. 2019 for constructing the lower bound with context (derivation in Appendix A). The expectations in this objective can be optimized with gradient ascent on samples drawn from the encoder using the reparameterization trick (Kingma & Welling (2014)).
@@ 4.1.3 Actor-Critic Policy
We largely follow Dreamer(v3) regarding training an actor-critic policy on imagined trajectories. However, we introduce context into policy learning by conditioning the actor and critic networks with the context. The actor is optimized to maximize the expected return on the imagined trajectories.
@@ 4.2 Naive Use of Context in Dreamer
As discussed previously, Dreamers ZSG capabilities have not been explored in an observable context setting, nor, to the best of our knowledge, has this setting been explored in MBRL in general. Thus, we propose and discuss naive learning variants from and within the contextual setting and contrast them with our proposed cRSSM. The naive variants are then used as baselines in our experiments.
@@ 4.2.1 Context as an Observation
A commonly adopted approach to incorporate the inferred or true context into an algorithm is to concatenate it with the state or observation (Perez et al., 2020; Biedenkapp et al., 2022; Sodhani et al., 2022; Benjamins et al., 2023). We study applying this approach to vanilla Dreamer. Vanilla Dreamer optimizes all the objectives defined in Section 4.1, but it does so without incorporating context in those objectives. To incorporate context naively, we provide it concatenated with the observation to the stochastic state encoder. where . Note that only the encoder gets the observation (or here the observation with context) as an input. The decoder then has to learn to reconstruct the context as it is part of the observation. The latent dynamics predictor used for imagination does not condition on observations. For a consistent imagination, the RSSM is burdened with retaining the context value (provided as an observation) which got encoded into the latent state, from which imagination begins. Since we only provide the latent state inferred by the encoder to the actor-critic model in this setting, if the context is not retained, then the actor-critic network will also not have access to the context information to learn accurate policies. This could make it hard to generalize OOD. Still, the simplicity of directly using context as part of the observation is appealing, which has led to it being the predominant approach in model-free cRL.
@@ 5 Experiments and Discussion
In this section, we assess the performance of Dreamer in achieving generalization under observed contexts. We compare various context conditioning approaches, following the evaluation protocols outlined in (Kirk et al., 2023). In particular, our findings highlight the effectiveness of our cRSSM method, showcasing quantitative and qualitative results in terms of zero-shot generalization (ZSG), particularly in scenarios involving changes to context parameters affecting the dynamics.
@@
Our experiments leverage CARL, tailored for our investigation into ZSG. In CARL we pick the following environments and contexts.
As these environments are good examples of the desired continuous contextual settings, they provide a suitable benchmark for our study. CartPole serves as a simple problem, while DMC Walker presents a more complex challenge. This approach aligns with prior works, such as Sodhani et al. (2022) and Zhang et al. (2021a), which evaluate generalization on DMControl tasks to changes in context.
@@
Single context variation: We sample context values uniformly for one context dimension in its training range, keeping the other fixed to its default value and vice versa.
@@
In the evaluation protocol, each context could constitute a task of different difficulty. For example, learning to control an agent in a context where it carries lighter loads might be easier than the one with heavier loads. Following Benjamins et al. (2023), to obtain an upper bound of the policy returns, we train expert agents for selected contexts that broadly cover our ranges of training and evaluation contexts. Expert agent performances are the best mean return over episodes among five seeds. This gives an approximate upper bound on the returns achievable if Dreamer is trained in a particular context. We used the best random policy mean return on episodes over five seeds to define a lower bound. See Section C.1 for the detailed expert and random agents performance.
@@
Setting | I | E | I+E | I | E | I+E | I | E | I+E | I | E | I+E |
CartPole | Walker | |||||||||||
Featurized | Pixel | |||||||||||
( d) | 1.000 | 1.000 | - | 1.000 | 0.938 | - | 0.903 | 0.561 | - | 0.940 | 0.546 | - |
( h) | 1.000 | 1.000 | - | 1.000 | 0.995 | - | 0.967 | 0.764 | - | 0.945 | 0.708 | - |
( c) | 1.000 | 1.000 | - | 1.000 | 0.997 | - | 0.966 | 0.769 | - | 0.966 | 0.733 | - |
( cR) | 1.000 | 1.000 | - | 1.000 | 1.000 | - | 0.985 | 0.806 | - | 0.959 | 0.710 | - |
( d) | 1.000 | 0.995 | - | 0.677 | 0.059 | - | 0.885 | 0.479 | - | 0.959 | 0.461 | - |
( h) | 1.000 | 0.996 | - | 1.000 | 0.169 | - | 0.959 | 0.571 | - | 0.947 | 0.571 | - |
( c) | 1.000 | 0.987 | - | 1.000 | 0.210 | - | 0.926 | 0.597 | - | 0.983 | 0.635 | - |
( cR) | 1.000 | 1.000 | - | 1.000 | 0.374 | - | 0.998 | 0.674 | - | 0.994 | 0.623 | - |
( d) | 1.000 | 0.945 | 0.998 | 0.901 | 0.038 | 0.210 | 0.842 | 0.520 | 0.595 | 0.915 | 0.503 | 0.570 |
( h) | 1.000 | 0.989 | 1.000 | 1.000 | 0.149 | 0.701 | 0.966 | 0.764 | 0.843 | 0.952 | 0.666 | 0.772 |
( c) | 1.000 | 0.970 | 1.000 | 1.000 | 0.257 | 0.779 | 0.972 | 0.727 | 0.830 | 0.965 | 0.724 | 0.823 |
( cR) | 1.000 | 0.997 | 1.000 | 1.000 | 0.334 | 0.826 | 0.982 | 0.677 | 0.820 | 0.988 | 0.691 | 0.843 |
@@ 5.1.1 How Effective is Domain Randomization for Dreamer’s ZSG?
To answer this, we compare the two approaches default-context, with the agent trained on the default context , and the hidden-context, which involves training with domain randomization of contexts.
@@
In summary, domain randomization benefits the ZSG of the Dreamer algorithm, and the improvement is striking for the pronounced pixel modality on the evaluated tasks.
@@ 5.1.2 Does Explicit Context Conditioning Aid ZSG?
Having established the benefits of domain randomization through hidden-context for Dreamer’s ZSG on the given tasks, our focus shifts to evaluating the impact of explicit context conditioning methods, namely cRSSM, our principled way to incorporate context into Dreamer’s world model; and concat-context where we augment the observations with the context.
@@ 5.1.3 Beyond Task Performance, Does Context-Conditioning Impact the Latent States?
To evaluate how the methods use context, we visually investigate Dreamer’s reconstruction of visually observable OOD contexts, which are either encoded into the latent state in concat-context or conditioned separately cRSSM. We choose the visually observable context to be the pole length in Cartpole, evaluating lengths shorter and longer than what it has seen in the training distribution (labeled short and long in Figure 5). A model capable of generating images conditioned on novel context demonstrates a semantic understanding of the context that is well grounded in the image space. In cases involving context-conditioned models, we also provide counterfactual visual explanations, exploring how the reconstructed pixel observation inferred from the original observation that encodes the factual context () differs if conditioned on a counterfactual context ().
Context: short
Context: long
Context: long
Context: short
@@
Our investigation reveals clear evidence of extrapolation capabilities in our proposed principled cRSSM approach compared to the vanilla concat-context strategy. Furthermore, through our visual counterfactual explanations, we observe indications that the latent state in cRSSM appears disentangled from the context, which explains the observed gains in generalization using this approach.
@@ 6 Conclusion and Future Work
We studied zero-shot generalization in Dreamer-style model-based reinforcement learning through the lens of contextual reinforcement learning. We discussed naive ways to incorporate contextual information into the MBRL learning pipeline and formulated the novel cRSSM for Dreamer. Our cRSSM provides a systematic approach to using context in the world modelling objectives under partial observability. Our experiments, using a rigorous evaluation protocol for zero-shot generalization, showed that naive approaches, such as domain randomization improve generalization performance. However, more principled methods such as our cRSSM are required to perform significantly better in-distribution and out-of-distribution. Our study opens the door to future work on zero-shot generalization for MBRL approaches such as Dreamer. Creating contextual benchmarks for environments such as Atari (Bellemare et al., 2013), DMLab (Beattie et al., 2016), ProcGen (Anand et al., 2021), and Minecraft (Guss et al., 2019) would be an interesting avenue for future research into ZSG, as contextual changes in these benchmarks would necessitate more pronounced changes in policies. We discuss this in more detail in Appendix G.